# Magnetic Dipole Portal Vector Dark Matter at Fixed-Targets

### v1

@@ 5.4 Visible Searches
Finally, the $Z^{\prime}$ decays visibly to SM final states $f^{+}f^{-}$, the only decay channel kinematically allowed, and can be searched for as displaced visible energy depositions in the detectors Berlin:2018bsc ; Andreas:2013lya ; Gninenko:2013rka ; LDMX:2025pkp or as missing energy/momentum if it decays outside of the detector. We consider displaced visible energy deposition limits from existing experiments NA64 NA64:2019auh , electron beam dump experiments E774, E141, Orsay, KEK, and E137 Andreas:2012mt , proton beam dumps CHARM and NuCal Blumlein:2011mv ; Tsai:2019buq .
For LDMX Phase II projections, we compute limits from displaced visible events and also events from missing energy/momentum in which the $Z^{\prime}$ decays outside of the detector. The number of events from a visibly decaying $Z^{\prime}$ is Berlin:2018bsc ,
@@
For LDMX Phase II we take $10^{16}$ EOT, with $z_{\text{min}}=43$ cm and $z_{\text{max}}=315$ cm for the visible signature, and $z_{\text{min}}=315$ cm and $z_{\text{max}}=\infty$ for the invisible signature. We use the approximate expression for the number of $Z^{\prime}$s produced in beam-target collisions as a function of $m_{Z^{\prime}}$ and couplings Berlin:2018bsc ,
@@ 6 Relic Targets, Constraints, and LDMX Projections
Figure 7: Summary of the most relevant constraints on this model in the $\sin\zeta$ vs $m_{\text{DM}}$ parameter space, overlaid with the relic targets (curves reproducing the observed relic abundance consistent with Planck) Planck:2018vyg shown as thick solid lines. The three subplots differ by the value of the $g_{D}$ parameter: with $g_{D}=10^{-4}$ on the left, $g_{D}=10^{-3}$ in the middle, and $g_{D}=10^{-2}$ on the right. 90% confidence level projections for LDMX Phase II LDMX:2025pkp are drawn in red: invisible vector meson decay is drawn in dashed, and visible decay of the $Z^{\prime}\to e^{+}e^{-}$ outside of the detector in dash-dot and inside of the detector in dotted. 90% confidence levels for NA64 NA64:2025ddk from invisible vector meson decay and visible $Z^{\prime}$ decay NA64:2018lsq at current statistics are in blue. The pink curves indicate regions in which the parameters are constrained by CMB measurements Slatyer:2015jla at 95% confidence level. Direct detection 90% confidence level constraints from DAMIC-M are plotted in green DAMIC-M:2025luv and from PANDAX-4T in orange PandaX:2024qfu . Grey regions are excluded at 95% confidence level by the beam dump experiments CHARM, E774, E141, Orsay, KEK, and E137 Andreas:2012mt , and at 90% by NuCal Tsai:2019buq , along with collider experiments LHCb LHCb:2019vmc , KLOE KLOE-2:2018kqf , BaBar BaBar:2014zli , NA48/2 NA482:2015wmo , and FASER FASER:2023tle at 90%.
@@
Constraints from the null result of existing beam dump and collider experiments searching for visible signatures from dark sectors, indicated by grey shaded regions, provide large coverage over the parameter space. Interestingly, for both LDMX and NA64, the sensitivity from invisible meson decays surpasses that of dark bremsstrahlung (including dark Higgs-strahlung), as discussed in more detail in Section 5. Furthermore, direct detection from DAMIC-M, in green, provide even stronger constraints. Limits from the CMB extend the constrained parameter region for $g_{D}=10^{-2}$, where they slightly outperform existing collider constraints. Although projections of LDMX Phase II (in red), for both visible and invisible searches, do not dramatically extend the sensitivity, it will however provide valuable validation of previous results, such as those from FASER. These results are not unexpected; being in the kinematic regime where the mediator cannot be produced on-shell leads to off-shell suppression of invisible final states, and the availability of visible search channels. Notice how all constraints that are dependent on the dark sector coupling, $g_{D}$, (direct detection, CMB, and invisible vector meson decays at fixed targets), become stronger compared to visible searches, as $g_{D}$ becomes larger. These results highlight the complementarity between different types of DM searches – from cosmological, laboratory, and astrophysical, searches.
@@ 7 Conclusion
We extend our previous work Catena:2023use on light spin-1 DM at LDMX (and other relevant fixed target experiments) to the case with a reversed mass hierarchy, $m_{Z^{\prime}}<2m_{\text{DM}}$, which alters the phenomenology and thus provides new search opportunities at LDMX and other relevant experiments. We demonstrate a viable light spin-1 DM model which includes an extension to the SM gauge symmetry by a new $SU(2)$, a $Z^{\prime}$ mediator, two new scalars, and a higher dimensional portal interaction between the dark sector and SM. The extended scalar sector that we outline, is capable of explaining the mass generation of DM and the $Z^{\prime}$, and also plays a role phenomenologically at fixed target experiments.
The relic abundance of DM is computed, where the DM undergoes freeze-out in the early universe. The annihilation processes dominantly responsible for setting the relic abundance range from dark sector only processes in the forbidden regime, to direct annihilation into SM fermions at larger portal couplings.
Our primary focus is on processes giving rise to signatures at fixed target experiments LDMX and NA64, under this reversed mass hierarchy scenario. While on-shell invisible mediator production is kinematically forbidden, we consider: off-shell dark bremsstrahlung into DM final states, dark Higgs-strahlung into dark scalar final states, invisible vector meson decays, and visible searches through decays of the $Z^{\prime}$ mediator to SM. We compute complementary constraints from direct detection, CMB, collider searches, and BBN. The dominant constraints come from DAMIC-M and PANDAX-4T via DM-electron scattering, collider and beam dump visible searches, and the CMB. A region of the parameter space remains unconstrained at $g_{D}\sim 10^{-2}$, $m_{\text{DM}}\gtrsim 100$ GeV.
@@ Acknowledgements.
Appendix A Invisible Vector Meson Decay Rate
@@
where $g_{V}=\frac{e\big(4\sin^{2}\theta\sin\rho-\sin\rho+3\cos\rho\tan\zeta\sin\theta_{w}\big)}{4\cos\theta_{w}\sin\theta_{w}}$, where $f_{V}$ is the form factor for the coupling of mesons with a vector current taken from Table III of Schuster:2021mlr which is derived from Bharucha:2015bzk .

### v2

@@ 5.4 Visible Searches
Finally, the $Z^{\prime}$ decays visibly to SM final states $f^{+}f^{-}$, the only decay channel kinematically allowed, and can be searched for as displaced visible energy depositions in the detectors Berlin et al. (2019); Andreas and others (2013); Gninenko (2014); Appert and others (2025) or as missing energy/momentum if it decays outside of the detector. We consider displaced visible energy deposition limits from existing experiments NA64 Banerjee and others (2020), electron beam dump experiments E774, E141, Orsay, KEK, and E137 Andreas et al. (2012), proton beam dumps CHARM and NuCal Blumlein and Brunner (2011); Tsai et al. (2021).
For LDMX Phase II projections, we compute limits from displaced visible events and also events from missing energy/momentum in which the $Z^{\prime}$ decays outside of the detector. The number of events from a visibly decaying $Z^{\prime}$ is Berlin et al. (2019),
@@
For LDMX Phase II we take $10^{16}$ EOT, with $z_{\text{min}}=43$ cm and $z_{\text{max}}=315$ cm for the visible signature, and $z_{\text{min}}=315$ cm and $z_{\text{max}}=\infty$ for the invisible signature. We use the approximate expression for the number of $Z^{\prime}$s produced in beam-target collisions as a function of $m_{Z^{\prime}}$ and couplings Berlin et al. (2019),
@@
6 Other Relevant Bounds
6.1 Colliders: Direct Search, Electroweak Precision, and Higgs Constraints
$Z^{\prime}$s decaying into SM final states, either promptly or long lived, after being produced in proton-proton or $e^{+}e^{-}$ collisions are searched for at collider experiments. We consider the limits from the null results of the experiments LHCb Aaij and others (2020), KLOE Anastasi and others (2018), BaBar Lees and others (2014), NA48/2 Batley and others (2015), and FASER Abreu and others (2024), most relevant for the sub-GeV DM mass range.
New physics is capable of affecting SM electroweak precision observables, such as neutral current interactions, leading to constraints on the parameter space of the beyond the SM model. Following Burgess et al. (1994); Tran and Yuan (2025); Alonso-González et al. (2025), we determine the effect of this model as a function of $\sin\zeta$ on the electroweak oblique parameters $S$ and $T$ Peskin and Takeuchi (1992). The neutral current interaction in our scenario can be split into the SM and the beyond the SM contributions,
$\mathcal{L}_{\text{nc}}=\mathcal{L}_{\text{SM,nc}}-\frac{e\sin\rho\tan\zeta\sin\theta_{w}}{\sin\theta_{w}\cos\theta_{w}}\sum_{i}\bar{f_{i}}\gamma^{\mu}\Big(-T_{3i}P_{L}+Q_{i}\Big)f_{i}Z_{\mu},$ (16)
where $Q_{i}$ is the charge of the corresponding fermion and $T_{3i}$ is the fermion’s third component of weak isospin. We match the above equation to equation 23 of Burgess et al. (1994) in order to write expressions for the $S$ and $T$ parameters in our model. Using the experimentally determined values from Navas and others (2024), $S=-0.05\pm 0.07$ and $T=0.00\pm 0.06$, we compute the $\chi^{2}$ as outlined in Alonso-González et al. (2025) and find the following 95% confidence level (C.L.) upper limit,
$\sin\zeta\leq 0.02.$ (17)
We find that these constraints are not competitive with the other constraints considered in this study. Furthermore, the predicted shifts in the $Z$-boson mass are sufficiently small that they do not impose additional meaningful bounds on the parameter space. We leave a more comprehensive analysis of the loop-induced electroweak contributions arising from the extended scalar sector, which may further modify the $S$ and $T$ parameters in addition to the Z-boson mass, to future work.
The invisible branching ratio of Higgs decays are excluded above 0.145 at 95% C.L. Aad and others (2022). In this scenario, the Higgs can decay invisibly through the process $H\to Z^{\prime}Z^{\prime}$, with a decay rate given in Eq. 56. The branching ratio of this process is small for parameter values that we consider, with a value of $\text{BR}_{H\to Z^{\prime}Z^{\prime}}\approx 4.6\times 10^{-13}$ for $\sin\zeta=0.1$ and $m_{\text{DM}}=1$ GeV. Therefore, we remain safe from invisible Higgs decay constraints.
6.2 Astrophysical and Cosmological Constraints
Energy Injection into the CMB
The annihilation of DM particles in the early universe injects energy into the CMB, which is constrained by current observations since it can change the recombination history, thus modifying the measured temperature and polarization power spectra Ade and others (2016). For constraints from the CMB to be relevant, DM must annihilate with a cross section that is s-wave, such that it is not velocity suppressed [^5] There does indeed exist bounds on p-wave annihilating DM Liu et al. (2016), however they impose constraints that are not competitive with the other constraints we already consider in this work.. In this model, DM annihilation into photons through the diagrams displayed in Fig. 2 are s-wave dominant, therefore the cross section is constant at low temperatures. We thus apply the 95% C.L. bound derived in Slatyer (2016) on the channel $\text{DM DM}\to\gamma\gamma$.
Big Bang Nucleosynthesis
Dark sector particles, especially of sub-GeV mass, thermally coupled to the SM and present during Big Bang Nucleosynthesis (BBN) can lead to alterations in the number of relativistic degrees of freedom, $N_{eff}$, affecting primordial abundances of light elements. The abundances of light elements are measured to a given precision, and any expected deviations from these observations lead to constraints on dark sector models Giovanetti et al. (2022); Sabti et al. (2020); Depta et al. (2019). A detailed consideration of constraints from BBN is beyond the scope of this work, however we include a conservative lower bound on the mass of the lightest dark species, $Z^{\prime}$, to be >10>10 MeV. Thanks to the large dark coupling values, $g_{D}=10^{-4},10^{-3},10^{-2}$, the $Z^{\prime}$ remains in thermal equilibrium with the SM thermal bath for all values of $\sin\zeta$ (down to $\sin\zeta=10^{-10}$), therefore $m_{Z^{\prime}}>10$ MeV applies across the entire parameter space considered here.
Indirect Detection
Searches of DM annihilation products in astrophysical environments constrain the annihilation cross section of DM Cirelli et al. (2025). DM annihilation into SM fermion final states such as $e^{+}e^{-}$, $q\bar{q}$, proceed via p-wave (velocity) suppressed cross sections, leaving constraints through these channels weak. We thus only consider the velocity independent annihilation into $\gamma\gamma$ channel, drawn in Fig. 2 (right), where the dominant constraints come from CMB as discussed above.
Bullet Cluster
DM self-interactions are constrained by measurements of the colliding galaxy cluster, the Bullet Cluster Markevitch et al. (2004); Wittman et al. (2018). The cross section of DM self-interactions at non-relativistic energies, for all parameter values considered in this study, fall safely below the upper limit of $\sigma/m_{\text{DM}}<1\>\text{cm}/\text{g}$ reported in Markevitch et al. (2004).
@@ 7 Relic Targets, Constraints, and LDMX Projections
Figure 7: Summary of the most relevant constraints on this model in the $\sin\zeta$ vs $m_{\text{DM}}$ parameter space, overlaid with the relic targets (curves reproducing the observed relic abundance consistent with Planck) Aghanim and others (2020) shown as thick solid lines. The three subplots differ by the value of the $g_{D}$ parameter: with $g_{D}=10^{-4}$ on the left, $g_{D}=10^{-3}$ in the middle, and $g_{D}=10^{-2}$ on the right. 90% C.L. projections for LDMX Phase II Appert and others (2025) are drawn in red: invisible vector meson decay is drawn in dashed, and visible decay of the $Z^{\prime}\to e^{+}e^{-}$ outside of the detector in dash-dot and inside of the detector in dotted. 90% C.L.s for NA64 Andreev and others (2025) from invisible vector meson decay and visible $Z^{\prime}$ decay Banerjee and others (2018) at current statistics are in blue. The pink curves indicate regions in which the parameters are constrained by CMB measurements Slatyer (2016) at 95% C.L.. Direct detection 90% C.L. constraints from DAMIC-M are plotted in green Aggarwal and others (2025) and from PANDAX-4T in orange Bo and others (2025). Grey regions are excluded at 95% C.L. by the beam dump experiments CHARM, E774, E141, Orsay, KEK, and E137 Andreas et al. (2012), and at 90% by NuCal Tsai et al. (2021), along with collider experiments LHCb Aaij and others (2020), KLOE Anastasi and others (2018), BaBar Lees and others (2014), NA48/2 Batley and others (2015), and FASER Abreu and others (2024) at 90%.
@@
Constraints from the null result of existing beam dump and collider experiments searching for visible signatures from dark sectors, indicated by grey shaded regions, provide large coverage over the parameter space. Interestingly, for both LDMX and NA64, the sensitivity from invisible meson decays surpasses that of dark bremsstrahlung (including dark Higgs-strahlung), as discussed in more detail in Section 5. Limits from the CMB extend the constrained parameter region for $g_{D}=10^{-2}$, where they slightly outperform existing collider constraints. Although projections of LDMX Phase II (in red), for both visible and invisible searches, do not dramatically extend the sensitivity, it will however provide valuable validation of previous results, such as those from FASER. Furthermore, direct detection from DAMIC-M and PANDAX-4T, in green and orange respectively, provide even stronger constraints than what LDMX is expected to provide. These results are not unexpected; being in the kinematic regime where the mediator cannot be produced on-shell leads to off-shell suppression of invisible final states, and the availability of visible search channels. Notice how all constraints that are dependent on the dark sector coupling, $g_{D}$, (direct detection, CMB, and invisible vector meson decays at fixed-targets), become stronger compared to visible searches, as $g_{D}$ becomes larger. These results highlight the complementarity between different types of DM searches – from cosmological, laboratory, and astrophysical, searches.
@@ 8 Conclusion
We consider a viable light spin-1 DM candidate that interacts with the SM through a dimension-5 portal that gives rise to off-shell invisible and visible production at fixed-target experiments such as NA64 and the future LDMX. Specifically, we explore a model where the SM gauge group is augmented by a non-Abelian dark $SU(2)_{D}$ symmetry, which is spontaneously broken by the vacuum expectation values of a scalar doublet and triplet. The extended scalar sector that we outline, is capable of explaining the mass generation of DM and the $Z^{\prime}$, and also plays a role phenomenologically at fixed-target experiments. This framework naturally realizes an inverse mass hierarchy where the DM is heavier than the $Z^{\prime}$.
This mass relationship leads to a distinctive phenomenological landscape. The mediator is kinematically forbidden from decaying invisibly into DM pairs, therefore DM production at fixed-target experiments giving rise to missing energy/momentum signatures must proceed through suppressed off-shell processes: dark off-shell bremsstrahlung and invisible vector meson decay. While these missing energy/momentum signatures are usually the dominant probe in on-shell scenarios, they fall short of other probes such as direct detection in off-shell scenarios. In this work, we quantified the interplay between these different frontiers, demonstrating that direct detection provides a dominant probe in this regime where fixed-target signatures are suppressed by off-shell kinematics. Direct detection experiments DAMIC-M and PANDAX-4T are found to be the current dominant constraints for this model, providing critical sensitivity that complements and often surpasses the reach of traditional fixed-target and collider searches.
The relic abundance of DM is computed, where DM undergoes freeze-out in the early universe, and relic targets are drawn. The annihilation processes dominantly responsible for setting the relic abundance range from dark sector only processes in the forbidden regime, to direct annihilation into SM fermions at larger portal coupling. These relic targets are compared with the current relevant constraints coming from astrophysical, laboratory, and cosmological probes.
Our primary focus is on processes giving rise to signatures at fixed-target experiments LDMX and NA64, under this reversed mass hierarchy scenario. While on-shell invisible mediator production is kinematically forbidden, we consider: off-shell dark bremsstrahlung into DM final states, dark Higgs-strahlung into dark scalar final states, invisible vector meson decays, and visible searches through decays of the $Z^{\prime}$ mediator to SM. We compute complementary constraints from direct detection, CMB, collider searches, and BBN. The dominant constraints come from DAMIC-M and PANDAX-4T via DM-electron scattering, collider and beam dump visible searches, and the CMB. A region of the parameter space remains unconstrained at $g_{D}\sim 10^{-2}$, $m_{\text{DM}}\gtrsim 100$ GeV.
Beyond the specific bounds of this $SU(2)_{D}$ model, this work highlights a broader methodological necessity in the search for light DM: the requirement to move beyond the standard on-shell mediator paradigm. By rigorously quantifying the suppression of fixed-target signatures in the inverse mass hierarchy, we demonstrate that the perceived hierarchy of sensitivity between experimental frontiers is highly model-dependent. Our findings show that direct detection is an essential primary probe that covers critical blind spots left by accelerator experiments when DM production is off-shell. This study provides a template for future dark sector explorations, underscoring that a truly robust search strategy for sub-GeV DM must treat direct detection and fixed-target probes as equal partners.
@@ Acknowledgements.
Appendix A Mass Matrices and Mixing Angles
In this appendix we present the mass matrices and mxing angles in the gauge- and scalar-sector of the model. After electroweak and dark sector symmetry breaking, the neutral gauge bosons $B_{\mu}$, $W^{3}_{\mu}$, and $X^{0}_{\mu}$ mix to form the physical photon $A_{\mu}$, the SM-like $Z$ boson, and a new massive neutral state $Z^{\prime}_{\mu}$. The relation between gauge and mass eigenstates is given by,
$\left(\begin{array}[]{cc}B_{\mu}\\ W^{3}_{\mu}\\ X^{0}_{\mu}\end{array}\right)=\left(\begin{array}[]{ccc}\cos\theta_{w}&-\sin\theta_{w}\cos\rho+\tan\zeta\sin\rho&-\sin\theta_{w}\sin\rho-\tan\zeta\cos\rho\\ \sin\theta_{w}&\cos\theta_{w}\cos\rho&\cos\theta_{w}\sin\rho\\ 0&-\sin\rho/\cos\zeta&\cos\rho/\cos\zeta\end{array}\right)\left(\begin{array}[]{cc}A_{\mu}\\ Z_{\mu}\\ Z^{\prime}_{\mu}\end{array}\right).$ (27)
The masses of the physical neutral gauge bosons are,
$m_{Z}^{2}=m_{W}^{2}A_{Z}+m_{X^{0}}^{2}\sec^{2}\zeta\sin^{2}\rho\,,\quad m_{Z^{\prime}}^{2}=m_{W}^{2}A_{Z^{\prime}}+m_{X^{0}}^{2}\sec^{2}\zeta\cos^{2}\rho\,$, (28)
where the dimensionless coefficients $A_{Z}$ and $A_{Z^{\prime}}$ are given by,
$A_{Z}$ $=\Big[\sec^{2}\theta_{w}\cos^{2}\rho+\tan\theta_{w}\tan\zeta\big(\tan\theta_{w}\tan\zeta\sin^{2}\rho-\sec\theta_{w}\sin 2\rho\big)\Big],$
$A_{Z^{\prime}}$ $=\Big[\sec^{2}\theta_{w}\sin^{2}\rho+\tan\theta_{w}\tan\zeta\big(\tan\theta_{w}\tan\zeta\cos^{2}\rho+\sec\theta_{w}\sin 2\rho\big)\Big],$ (29)
and the parameter $\rho$ is defined as,
$\tan 2\rho\equiv\frac{m_{Z}^{2}\sin\theta_{w}\sin 2\zeta}{m_{\text{DM}}^{2}\cos^{2}\beta-m_{Z}^{2}(\cos^{2}\zeta-\sin^{2}\theta_{w}\sin^{2}\zeta)}.$ (30)
We now turn to the scalar sector. After the symmetry breaking in the dark sector, the quadratic terms for the dark charged and CP-even neutral scalars are given by,
$\mathcal{L}_{\rm mass}=\frac{1}{2}\left(\phi^{0}_{D}~\Sigma^{0}_{D}\right)\mathcal{M}_{0}^{2}\left(\begin{array}[]{c}\phi^{0}_{D}\\ \Sigma^{0}_{D}\end{array}\right)+\left(\phi^{+}_{D}~\Sigma^{+}_{D}\right)\mathcal{M}_{\pm}^{2}\left(\begin{array}[]{c}\phi^{-}_{D}\\ \Sigma^{-}_{D}\end{array}\right)\,$, (35)
where,
$\mathcal{M}_{0}^{2}=\left(\begin{array}[]{cc}2\lambda_{\Phi_{D}}v_{\Phi}^{2}&\frac{v_{\Phi}}{2}(\sqrt{2}\lambda_{\Phi_{D}\Sigma_{D}}v_{\Sigma}-\mu_{\Phi_{D}\Sigma_{D}})\\ \frac{v_{\Phi}}{2}(\sqrt{2}\lambda_{\Phi_{D}\Sigma_{D}}v_{\Sigma}-\mu_{\Phi_{D}\Sigma_{D}})&\lambda_{\Sigma_{D}}v_{\Sigma}^{2}+\frac{\mu_{\Phi_{D}\Sigma_{D}}v_{\Phi}^{2}}{2\sqrt{2}v_{\Sigma}}\end{array}\right)\,\quad,\mathcal{M}_{\pm}^{2}=\mu_{\Phi_{D}\Sigma_{D}}\left(\begin{array}[]{cc}v_{\Sigma}/\sqrt{2}&v_{\Phi}/2\\ v_{\Phi}/2&v_{\Phi}^{2}/2\sqrt{2}v_{\Sigma}\end{array}\right).$ (40)
The mixing angles relating the gauge and mass eigenstates are given by the following relations,
$\left(\begin{array}[]{cc}\phi^{\pm}_{D}\\ \Sigma^{\pm}_{D}\end{array}\right)=\left(\begin{array}[]{cc}\cos\beta&-\sin\beta\\ \sin\beta&\cos\beta\end{array}\right)\left(\begin{array}[]{cc}\hat{\phi}^{\pm}_{D}\\ \hat{\Sigma}^{\pm}_{D}\end{array}\right)\,\quad\left(\begin{array}[]{cc}\phi^{0}_{D}\\ \Sigma^{0}_{D}\end{array}\right)=\left(\begin{array}[]{cc}\cos\theta_{t}&-\sin\theta_{t}\\ \sin\theta_{t}&\cos\theta_{t}\end{array}\right)\left(\begin{array}[]{cc}\hat{\phi}^{0}_{D}\\ \hat{\Sigma}^{0}_{D}\end{array}\right),$ (53)
where,
$\tan\beta=\frac{v_{\Phi}}{\sqrt{2}v_{\Sigma}}\,,\quad\tan 2\theta_{t}=\frac{2(\mathcal{M}^{2}_{0})_{12}}{(\mathcal{M}^{2}_{0})_{22}-(\mathcal{M}^{2}_{0})_{11}}\,.$ (54)
Appendix B Decay Rates
@@
where $g_{V}=\frac{e\big(4\sin^{2}\theta\sin\rho-\sin\rho+3\cos\rho\tan\zeta\sin\theta_{w}\big)}{4\cos\theta_{w}\sin\theta_{w}}$, where $f_{V}$ is the form factor for the coupling of mesons with a vector current taken from Table III of Schuster et al. (2022) which is derived from Bharucha et al. (2016).
The invisible decay rate of the Higgs is given by,
[…]
$\times\frac{\Big(\cos^{2}\theta_{w}\sin\rho+\sin\theta_{w}(\sin\rho\sin\theta_{w}+\cos\rho\tan\zeta)\Big)^{4}v_{H}^{2}}{512\pi\cos^{4}\theta_{w}m_{Z^{\prime}}^{4}\sin^{2}\theta_{w}m_{H}^{3}},$
where $v_{H}$ is the Higgs vacuum expectation value.
Appendix C Feynman Rules
Here we present the relevant feynman rules.
New Dark Gauge Bosons
$=$ $ig_{D}\Big[C_{1}\Big(\eta_{\mu\nu}\left(p^{\prime}_{\alpha}-p_{\alpha}\right)-\eta_{\alpha\nu}\left(2p^{\prime}_{\mu}+p_{\mu}\right)+\eta_{\mu\alpha}\left(p^{\prime}_{\nu}+2p_{\nu}\right)\Big)$ (57)
$+C_{2}\sin\zeta\Big(\eta_{\alpha\mu}q_{\nu}-\eta_{\alpha\nu}q_{\mu}\Big)\Big]$
$C_{1}=\begin{cases}0,&\gamma\\ \frac{\cos\rho}{\cos\zeta},&Z^{\prime}\\ \frac{\sin\rho}{\cos\zeta},&Z\end{cases}$ (58)
$C_{2}=\begin{cases}\cos\theta_{w}\,,&\gamma\\ \left(\sin\theta_{w}\cos\rho+\tan\zeta\sin\rho\right)\,,&Z\\ -\left(\sin\theta_{w}\sin\rho+\tan\zeta\cos\rho\right)\,,&Z^{\prime}\end{cases}$ (59)
$=$ $\>\frac{ie\gamma_{\mu}}{4\cos\theta_{w}\sin\theta_{w}}\bigg[\Big(\sin\rho(4\sin^{2}\theta_{w}-1)+3\cos\rho\tan\zeta\sin\theta_{w}\Big)$ (60)
$+\gamma_{5}\Big(\sin\rho+\cos\rho\tan\zeta\sin\theta_{w}\Big)\bigg]$
New Scalar Interactions
$\frac{ig_{D}\cos\beta\sin\zeta}{m_{\text{DM}}\sin\beta}\Big[C_{1,S}\eta_{\nu\alpha}-C_{2,S}q_{\nu}p^{\prime}_{\alpha}\Big]$ (61)
$C_{1,S}=\begin{cases}\cos\theta_{w}(p^{\prime}\cdot q),&\gamma\\ \frac{4m_{\text{DM}}^{2}\sin\rho\sin^{2}\beta}{\cos\zeta\sin\zeta}+(p^{\prime}\cdot q)\left(\tan\zeta\sin\rho-\sin\theta_{w}\cos\rho\right)\,,&Z\\ -\frac{4m_{\text{DM}}^{2}\sin\rho\sin^{2}\beta}{\cos\zeta\sin\zeta}-(p^{\prime}\cdot q)\left(\tan\zeta\cos\rho+\sin\theta_{w}\sin\rho\right)\,,&Z^{\prime}\end{cases}$ (62)
$C_{2,S}=\begin{cases}\cos\theta_{w},&\gamma\\ \left(\cos\rho\sin\theta_{w}-\sin\rho\tan\zeta\right)\,,&Z\\ \left(\sin\rho\sin\theta_{w}+\cos\rho\tan\zeta\right)\,,&Z^{\prime}\end{cases}$ (63)
[…]
$=\>ig_{D}\frac{\cos\rho}{\cos\zeta}\big(\cos^{2}\beta+\sin^{2}\beta/2\big)\big(p_{\alpha}-p^{\prime}_{\alpha}\big)$ (66)
[…]